# Computational Economics in Large Language Models: Exploring Model Behavior and Incentive Design under Resource Constraints

### v1

@@ 1 Introduction
The advent of Large Language Models (LLMs) has marked a pivotal moment in artificial intelligence, demonstrating remarkable capabilities that are reshaping both the scientific landscape and numerous industries [1, zhou2022contrastive]. These models, through unprecedented scaling of parameters, data, and computation, have exhibited emergent abilities [2] that were not explicitly programmed, [3].allowing them to perform complex reasoning tasks via techniques like chain-of-thought prompting [4] and even deliberate problem-solving through more sophisticated strategies [5][6]. The principles governing their performance gains, often described by predictable scaling laws [7], have fueled a race towards ever-larger models. This trend is further amplified by the extension of these architectures into the multi-modal domain, tackling complex tasks such as audio-visual event analysis and efficient video grounding [8], which inherently demand even greater computational resources.
@@
However, this extraordinary progress comes at a staggering and often prohibitive cost. The immense computational and energy requirements for training and deploying state-of-the-art models pose a significant barrier to their democratization, accessibility, and environmental sustainability [9].[10] This reliance on massive computational power has created a ”hardware lottery,” where the viability of a research idea can be determined as much by its compatibility with existing hardware as by its intrinsic merit [11]. As economists have noted, while AI has drastically reduced the cost of prediction, the associated judgment and infrastructure costs[12] remain substantial [13]. This economic reality necessitates a paradigm shift from a focus on pure performance to a more holistic consideration of performance under strict resource constraints.
@@
In response, the research community has explored numerous avenues for improving model efficiency. Architectural innovations, such as the sparsely-gated Mixture-of-Experts (MoE) layer [14], have[15] enabled the scaling to trillion-parameter models by activating only a fraction of the network for each input [16, 17]. This paradigm has been successfully implemented in powerful open-source models [18] and represents a major direction in efficient model design, as detailed in comprehensive surveys [19]. Concurrently, efforts to design more fundamentally efficient attention mechanisms have yielded architectures like the Reformer [20] and Longformer [21], which reduce the quadratic complexity of self-attention, enabling models to process much longer sequences [22]. Other approaches focus on dynamic computation, allowing models to adapt their computational depth based on input complexity, either through adaptive computation steps [23, 24] or dynamic early exiting from model layers [25].
@@
Complementary to architectural changes, model compression techniques aim to shrink dense models into more manageable forms. Seminal work on the lottery ticket hypothesis suggests that large networks contain sparse, highly trainable subnetworks that can be isolated [26]. This has inspired methods for structured pruning of tokens and heads [27]. Furthermore, knowledge distillation has proven to be a powerful technique for transferring the capabilities of a large ”teacher” model to a smaller ”student” model [28, 29, 30], a principle that has been extended to visual dialog systems [31] and even inspired progressive module replacement strategies [32]. Multi-objective convex quantization offers another path to compression by optimizing for multiple objectives simultaneously . This broad pursuit of efficiency is not unique to NLP. It is a central theme across AI, from developing robust, interference-aware wireless sensing systems for healthcare [33] and activity[gu2023mamba] recognition , to creating reliable facial expression recognition systems that can handle label noise and domain heterogeneity . The overarching goal remains the same: maximizing utility under constraints, whether they be computational, energetic, or related to data quality.
@@
Despite this wealth of techniques, a significant gap remains: the absence of a unified theoretical framework to understand and guide the internal resource allocation behavior of LLMs. While interpretability research has made strides in revealing what models learn, showing that they rediscover classical NLP pipelines [34] and that their feed-forward layers act as key-value memories [35], it often stops short of explaining why they behave as they do or how to steer this behavior. Indeed, it has been shown that simple interpretations, such as equating attention with explanation, can be misleading [36], and deeper grammatical analysis is required to understand what different components truly learn [37]. Simultaneously, we observe models exhibiting increasingly rational, agent-like behaviors, such as teaching themselves to use external tools [38], synergizing reasoning with action [39], [40]and engaging in self-collaboration to solve complex problems [41]. This emergent rationality suggests that an underlying, perhaps implicit, economic logic governs their operations, which current engineering-focused approaches do not fully capture. This is further highlighted by the need for models to handle open-set conditions, a challenge in fields like gesture recognition where systems must robustly manage uncertainty [42].
@@
This paper introduces the perspective of Computational Economics as a novel theoretical lens to address this gap. We propose to model an LLM not as a monolithic computational graph, but as an internal economic system composed of numerous, competing ”agents” (e.g., attention heads, neuron blocks) that must bid for and allocate finite computational resources to maximize a collective objective. This framework is grounded in established theories of algorithmic game theory [43] and the information bottleneck principle [44, 45], and it provides a principled foundation for designing and analyzing model behavior. By framing the problem in this way, we can leverage powerful concepts from mechanism design [46] to create explicit ”incentive structures”—for instance, through novel loss functions—that guide the model to learn more efficient and adaptive resource allocation strategies [47]. Such a principled approach has the potential to unify our understanding of efficiency and inform the development of more robust AI systems, from secure federated learning networks [48] to multi-modal systems for recognizing fine-grained human actions or emotions from diverse signals . It also forces us to consider the costs and risks associated with system design, a critical aspect in security domains like preventing physical layer attacks or acoustic eavesdropping.
@@ 2.1 Efficiency in Large Language Models
The pursuit of computational efficiency in LLMs has predominantly followed two paths: architectural innovation and model compression. Architectural innovations aim to fundamentally reduce the computational complexity of the Transformer architecture. The most prominent among these is the Mixture-of-Experts (MoE) paradigm, first proposed in early machine learning [14] and later scaled to create trillion-parameter yet computationally feasible models [16, 17]. The core idea is conditional computation, where only a sparse subset of ”expert” sub-networks is activated for any given input, a concept now central to leading open-source models [18] and extensively reviewed in recent surveys [19]. Another critical bottleneck is the quadratic complexity of the self-attention mechanism. To address this, researchers have developed more efficient attention variants, such as those employing locality-sensitive hashing [20] or combining local and global attention patterns [21], thereby extending the feasible context length of models dramatically [22]. The design of efficient network backbones is a shared goal across deep learning, with similar principles being applied to create unified static and dynamic networks for efficient video processing [8].
@@
Model compression techniques, on the other hand, seek to reduce the size and computational cost of pre-existing dense models. Pruning, inspired by seminal findings like the Lottery Ticket Hypothesis [26], involves removing redundant weights or structured components like entire attention heads [27]. Knowledge distillation offers a different approach, training a smaller ”student” model to mimic the output behavior of a larger ”teacher” model [28, 30]. This principle of transferring knowledge from a complex system to a simpler one has found broad applicability, for instance in developing context-aware visual dialog systems [31]. Quantization further reduces model size by representing weights with lower-precision data types, a process that can be framed as a multi-objective optimization problem to balance size and accuracy . These efficiency-driven efforts are not isolated to mainstream NLP and vision; they mirror challenges in specialized domains like creating anti-interference activity recognition systems from WiFi signals, which requires careful subcarrier selection to manage signal complexity and cost .
@@
Finally, dynamic computation methods allow a model’s computational budget to vary per input. This includes adaptive computation time in recurrent networks [23] and, more relevant to Transformers, early exiting strategies where ”easy” inputs are processed by fewer layers [25]. This concept of dynamic resource allocation based on task difficulty is a direct precursor to our economic framework. The challenge of creating robust systems that perform well under heterogeneous conditions is universal, whether in dynamic facial expression recognition or in federated learning across diverse edge networks .
@@ 2.2 Interpretability and Internal Mechanisms
While efficiency research focuses on how to make models cheaper, interpretability research asks what these models are actually learning. A significant body of work has sought to peer inside the ”black box.” Early studies revealed that deep language models like BERT implicitly learn a hierarchy of linguistic properties, effectively rediscovering the classical NLP pipeline from part-of-speech tagging to semantic roles across their layers [34]. Probing techniques have been instrumental in these analyses, training simple classifiers on a model’s internal representations to test for specific encoded knowledge [49, 50]. Other work has demystified specific components, for example, by showing that Transformer feed-forward layers function as distributed key-value memories [35].
The attention mechanism, once thought to be a straightforward window into a model’s reasoning, has been the subject of intense scrutiny. Foundational work has cautioned that high attention weights do not necessarily equate to explanatory importance [36], prompting more nuanced and rigorous methods for analysis, such as examining the grammatical roles learned by different attention heads [37]. Understanding these internal representations is crucial for building reliable systems. For instance, in affective computing, achieving robust emotion recognition requires fusing information from multiple modalities like vision and WiFi signals, and understanding how the model weighs each source is key [51]. Similarly, building systems that can suppress label noise in real-world data requires a model of the underlying generative process of errors .
More recently, research has focused on the emergent, agent-like behaviors of LLMs. Models can now learn to use external APIs and tools to augment their capabilities [38], create explicit reasoning steps before acting [39], and even form multi-agent conversational structures to collaboratively solve problems [52]. This emergent rationality, where models appear to make strategic choices, motivates our central question: can these behaviors be explained and guided by economic principles? The need for such principled guidance is evident in security-critical applications, where system vulnerabilities can be exploited, such as in fingerprint-based authentication or through side-channel attacks that eavesdrop on keystrokes .
@@ 2.3 Algorithmic Mechanism Design and Agent-Based Modeling
Our work draws its core theoretical inspiration from algorithmic game theory [43], particularly the subfield of mechanism design. Mechanism design is essentially the ”reverse engineering” of game theory; it focuses on designing the rules of a game to incentivize self-interested agents to behave in a way that achieves a desirable system-wide outcome [46]. This framework has been used to design systems that elicit truthful information from participants [53] and to develop contracts that incentivize effort in machine learning contexts [47]. Our key insight is to apply this thinking not to external, human agents, but to the internal components of a neural network itself.
This perspective is supported by the information bottleneck principle, which posits that any learning system should optimally trade off between compressing its input and preserving information relevant to its output [44], a concept that has been operationalized for deep neural networks [45]. This provides a formal language for reasoning about the trade-offs inherent in resource-constrained computation. Furthermore, the idea of treating system components as agents has parallels in other areas of AI. For example, research in federated learning explores how to orchestrate model migration and architecture search among heterogeneous edge devices, treating each device as a self-contained agent in a larger network [48].
Our framework also connects to the concept of algorithmic recourse, which studies how to advise individuals on changing their features to achieve a more favorable outcome from a model, explicitly modeling the ”cost” of change [54]. We transpose this idea inward, asking how the model itself can choose the most ”cost-effective” computational path. By viewing the model’s internal components as rational agents operating under scarcity, we can move beyond simply observing their behavior and begin to proactively design the economic incentives that govern their interactions. This is crucial for developing the next generation of AI systems, which must be not only powerful but also efficient, robust, and predictable, whether they are used for benchmarking micro-actions , providing in-home pulmonary function monitoring , or enabling robust open-set gesture recognition [42].
@@ 3.1 A Computational Economics Framework for LLMs
The foundational premise of our work is that the complex, multi-component architecture of an LLM can be productively analyzed as a microeconomic system. In this system, individual components act as rational agents that make local decisions to collectively optimize a global objective, all while operating under conditions of resource scarcity. This abstraction allows us to leverage the powerful analytical tools of economics and mechanism design [43].
@@ 3.1.1 Defining the Economic Agents and Resources
The primary computational resource we consider is selective attention and neural activation. In a standard Transformer, resources are allocated profligately; every token attends to every other token, and FFNs are typically dense. Our framework introduces the concept of a computational budget, $B$, which constrains the total amount of resources an agent or a layer can consume for a given input. This scarcity is the driving force of the economic behavior we aim to study. For example, the work on wide residual networks implicitly explores the trade-off in allocating parameter resources to width versus depth [55], while our focus is on the dynamic allocation of computational resources during inference.
@@ 3.1.2 Modeling Utility and Cost
Computational Cost: Every action taken by an agent incurs a computational cost, $C_{\text{comp}}$. This cost is a function of the resources consumed. For an attention head, the cost could be proportional to the number of tokens it strongly attends to. For an FFN, it could be the number of activated neurons. This aligns with the goals of dynamic computation methods [23, 25] but provides a more granular, agent-centric view. For a given Transformer layer $l$, we can define its computational cost as a function of its activation patterns. For instance, we can define the cost as the sum of the L1 norm of the attention scores and the FFN activations:
@@ 3.2.1 Implementing Resource Constraints
We implement resource constraints using techniques that induce sparsity in the model’s activations during inference. We focus on constraining the attention mechanism, as it is a primary driver of computational cost and has been a focus of efficiency research [20, 21]. We employ two main techniques:

### v2

@@ 1 Introduction
The advent of Large Language Models (LLMs) has marked a pivotal moment in artificial intelligence, demonstrating remarkable capabilities that are reshaping both the scientific landscape and numerous industries brown2020language. These models, through unprecedented scaling of parameters, data, and computation, have exhibited emergent abilities wei2022emergent that were not explicitly programmed, cui2025learningheterogeneitygeneralizingdynamic.allowing them to perform complex reasoning tasks via techniques like chain-of-thought prompting wei2022chain and even deliberate problem-solving through more sophisticated strategies yao2023tree. The principles governing their performance gains, often described by predictable scaling laws kaplan2020scaling, have fueled a race towards ever-larger models. This trend is further amplified by the extension of these architectures into the multi-modal domain, tackling complex tasks such as audio-visual event analysis and efficient video grounding , which inherently demand even greater computational resources.
@@
However, this extraordinary progress comes at a staggering and often prohibitive cost. The immense computational and energy requirements for training and deploying state-of-the-art models pose a significant barrier to their democratization, accessibility, and environmental sustainability bender2021dangers.9613773 This reliance on massive computational power has created a ”hardware lottery,” where the viability of a research idea can be determined as much by its compatibility with existing hardware as by its intrinsic merit hooker2020hardware. As economists have noted, while AI has drastically reduced the cost of prediction, the associated judgment and infrastructure costs10916982 remain substantial agrawal2022prediction. This economic reality necessitates a paradigm shift from a focus on pure performance to a more holistic consideration of performance under strict resource constraints.
@@
In response, the research community has explored numerous avenues for improving model efficiency. Architectural innovations, such as the sparsely-gated Mixture-of-Experts (MoE) layer shazeer2017outrageously, have10339891 enabled the scaling to trillion-parameter models by activating only a fraction of the network for each input lepikhin2020gshard, fedus2022switch. This paradigm has been successfully implemented in powerful open-source models jiang2024mixtral and represents a major direction in efficient model design, as detailed in comprehensive surveys fedus2022survey. Concurrently, efforts to design more fundamentally efficient attention mechanisms have yielded architectures like the Reformer kitaev2020reformer and Longformer beltagy2020longformer, which reduce the quadratic complexity of self-attention, enabling models to process much longer sequences ding2023longnet. Other approaches focus on dynamic computation, allowing models to adapt their computational depth based on input complexity, either through adaptive computation steps graves2016adaptive, 10508334 or dynamic early exiting from model layers xin2020deebert.
@@
Complementary to architectural changes, model compression techniques aim to shrink dense models into more manageable forms. Seminal work on the lottery ticket hypothesis suggests that large networks contain sparse, highly trainable subnetworks that can be isolated frankle2018the. This has inspired methods for structured pruning of tokens and heads wang2021spatten. Furthermore, knowledge distillation has proven to be a powerful technique for transferring the capabilities of a large ”teacher” model to a smaller ”student” model sanh2019distilbert, 9075376, jiao2019tinybert, a principle that has been extended to visual dialog systems and even inspired progressive module replacement strategies xu2020bert. Multi-objective convex quantization offers another path to compression by optimizing for multiple objectives simultaneously . This broad pursuit of efficiency is not unique to NLP. It is a central theme across AI, from developing robust, interference-aware wireless sensing systems for healthcare 10804189 and activity recognition , to creating reliable facial expression recognition systems that can handle label noise and domain heterogeneity . The overarching goal remains the same: maximizing utility under constraints, whether they be computational, energetic, or related to data quality.
@@
Despite this wealth of techniques, a significant gap remains: the absence of a unified theoretical framework to understand and guide the internal resource allocation behavior of LLMs. While interpretability research has made strides in revealing what models learn, showing that they rediscover classical NLP pipelines tenney2019bert and that their feed-forward layers act as key-value memories geva2021transformer, it often stops short of explaining why they behave as they do or how to steer this behavior. Indeed, it has been shown that simple interpretations, such as equating attention with explanation, can be misleading jain2019attention, and deeper grammatical analysis is required to understand what different components truly learn ghandi2023what. Simultaneously, we observe models exhibiting increasingly rational, agent-like behaviors, such as teaching themselves to use external tools schick2023toolformer, synergizing reasoning with action yao2022react, 10699348and engaging in self-collaboration to solve complex problems dong2023self. This emergent rationality suggests that an underlying, perhaps implicit, economic logic governs their operations, which current engineering-focused approaches do not fully capture. This is further highlighted by the need for models to handle open-set conditions, a challenge in fields like gesture recognition where systems must robustly manage uncertainty 10899398.
@@
This paper introduces the perspective of Computational Economics as a novel theoretical lens to address this gap. We propose to model an LLM not as a monolithic computational graph, but as an internal economic system composed of numerous, competing ”agents” (e.g., attention heads, neuron blocks) that must bid for and allocate finite computational resources to maximize a collective objective. This framework is grounded in established theories of algorithmic game theory nisan2007algorithmic and the information bottleneck principle tishby1999information, alemi2016deep, and it provides a principled foundation for designing and analyzing model behavior. By framing the problem in this way, we can leverage powerful concepts from mechanism design dutting2019machine to create explicit ”incentive structures”—for instance, through novel loss functions—that guide the model to learn more efficient and adaptive resource allocation strategies narayanan2020learning. Such a principled approach has the potential to unify our understanding of efficiency and inform the development of more robust AI systems, from secure federated learning networks 10251628 to multi-modal systems for recognizing fine-grained human actions or emotions from diverse signals . It also forces us to consider the costs and risks associated with system design, a critical aspect in security domains like preventing physical layer attacks or acoustic eavesdropping.
@@ 2.1 Efficiency in Large Language Models
The pursuit of computational efficiency in LLMs has predominantly followed two paths: architectural innovation and model compression. Architectural innovations aim to fundamentally reduce the computational complexity of the Transformer architecture. The most prominent among these is the Mixture-of-Experts (MoE) paradigm, first proposed in early machine learning shazeer2017outrageously and later scaled to create trillion-parameter yet computationally feasible models lepikhin2020gshard, fedus2022switch. The core idea is conditional computation, where only a sparse subset of ”expert” sub-networks is activated for any given input, a concept now central to leading open-source models jiang2024mixtral and extensively reviewed in recent surveys fedus2022survey. Another critical bottleneck is the quadratic complexity of the self-attention mechanism. To address this, researchers have developed more efficient attention variants, such as those employing locality-sensitive hashing kitaev2020reformer or combining local and global attention patterns beltagy2020longformer, thereby extending the feasible context length of models dramatically ding2023longnet. The design of efficient network backbones is a shared goal across deep learning, with similar principles being applied to create unified static and dynamic networks for efficient video processing .
@@
Model compression techniques, on the other hand, seek to reduce the size and computational cost of pre-existing dense models. Pruning, inspired by seminal findings like the Lottery Ticket Hypothesis frankle2018the, involves removing redundant weights or structured components like entire attention heads wang2021spatten. Knowledge distillation offers a different approach, training a smaller ”student” model to mimic the output behavior of a larger ”teacher” model sanh2019distilbert, jiao2019tinybert. This principle of transferring knowledge from a complex system to a simpler one has found broad applicability, for instance in developing context-aware visual dialog systems . Quantization further reduces model size by representing weights with lower-precision data types, a process that can be framed as a multi-objective optimization problem to balance size and accuracy . These efficiency-driven efforts are not isolated to mainstream NLP and vision; they mirror challenges in specialized domains like creating anti-interference activity recognition systems from WiFi signals, which requires careful subcarrier selection to manage signal complexity and cost .
@@
Finally, dynamic computation methods allow a model’s computational budget to vary per input. This includes adaptive computation time in recurrent networks graves2016adaptive and, more relevant to Transformers, early exiting strategies where ”easy” inputs are processed by fewer layers xin2020deebert. This concept of dynamic resource allocation based on task difficulty is a direct precursor to our economic framework. The challenge of creating robust systems that perform well under heterogeneous conditions is universal, whether in dynamic facial expression recognition or in federated learning across diverse edge networks .
@@ 2.2 Interpretability and Internal Mechanisms
While efficiency research focuses on how to make models cheaper, interpretability research asks what these models are actually learning. A significant body of work has sought to peer inside the ”black box.” Early studies revealed that deep language models like BERT implicitly learn a hierarchy of linguistic properties, effectively rediscovering the classical NLP pipeline from part-of-speech tagging to semantic roles across their layers tenney2019bert. Probing techniques have been instrumental in these analyses, training simple classifiers on a model’s internal representations to test for specific encoded knowledge niven2019probing. Other work has demystified specific components, for example, by showing that Transformer feed-forward layers function as distributed key-value memories geva2021transformer.
The attention mechanism, once thought to be a straightforward window into a model’s reasoning, has been the subject of intense scrutiny. Foundational work has cautioned that high attention weights do not necessarily equate to explanatory importance jain2019attention, prompting more nuanced and rigorous methods for analysis, such as examining the grammatical roles learned by different attention heads ghandi2023what. Understanding these internal representations is crucial for building reliable systems. For instance, in affective computing, achieving robust emotion recognition requires fusing information from multiple modalities like vision and WiFi signals, and understanding how the model weighs each source is key 10149418. Similarly, building systems that can suppress label noise in real-world data requires a model of the underlying generative process of errors .
More recently, research has focused on the emergent, agent-like behaviors of LLMs. Models can now learn to use external APIs and tools to augment their capabilities schick2023toolformer, create explicit reasoning steps before acting yao2022react, and even form multi-agent conversational structures to collaboratively solve problems wu2023autogen. This emergent rationality, where models appear to make strategic choices, motivates our central question: can these behaviors be explained and guided by economic principles? The need for such principled guidance is evident in security-critical applications, where system vulnerabilities can be exploited, such as in fingerprint-based authentication or through side-channel attacks that eavesdrop on keystrokes .
@@ 2.3 Algorithmic Mechanism Design and Agent-Based Modeling
Our work draws its core theoretical inspiration from algorithmic game theory nisan2007algorithmic, particularly the subfield of mechanism design. Mechanism design is essentially the ”reverse engineering” of game theory; it focuses on designing the rules of a game to incentivize self-interested agents to behave in a way that achieves a desirable system-wide outcome dutting2019machine. This framework has been used to design systems that elicit truthful information from participants evans2021truthful and to develop contracts that incentivize effort in machine learning contexts narayanan2020learning. Our key insight is to apply this thinking not to external, human agents, but to the internal components of a neural network itself.
This perspective is supported by the information bottleneck principle, which posits that any learning system should optimally trade off between compressing its input and preserving information relevant to its output tishby1999information, a concept that has been operationalized for deep neural networks alemi2016deep. This provides a formal language for reasoning about the trade-offs inherent in resource-constrained computation. Furthermore, the idea of treating system components as agents has parallels in other areas of AI. For example, research in federated learning explores how to orchestrate model migration and architecture search among heterogeneous edge devices, treating each device as a self-contained agent in a larger network 10251628.
Our framework also connects to the concept of algorithmic recourse, which studies how to advise individuals on changing their features to achieve a more favorable outcome from a model, explicitly modeling the ”cost” of change karimi2021survey. We transpose this idea inward, asking how the model itself can choose the most ”cost-effective” computational path. By viewing the model’s internal components as rational agents operating under scarcity, we can move beyond simply observing their behavior and begin to proactively design the economic incentives that govern their interactions. This is crucial for developing the next generation of AI systems, which must be not only powerful but also efficient, robust, and predictable, whether they are used for benchmarking micro-actions , providing in-home pulmonary function monitoring , or enabling robust open-set gesture recognition 10899398.
@@ 3.1 A Computational Economics Framework for LLMs
The foundational premise of our work is that the complex, multi-component architecture of an LLM can be productively analyzed as a microeconomic system. In this system, individual components act as rational agents that make local decisions to collectively optimize a global objective, all while operating under conditions of resource scarcity. This abstraction allows us to leverage the powerful analytical tools of economics and mechanism design nisan2007algorithmic.
@@ 3.1.1 Defining the Economic Agents and Resources
The primary computational resource we consider is selective attention and neural activation. In a standard Transformer, resources are allocated profligately; every token attends to every other token, and FFNs are typically dense. Our framework introduces the concept of a computational budget, $B$, which constrains the total amount of resources an agent or a layer can consume for a given input. This scarcity is the driving force of the economic behavior we aim to study. For example, the work on wide residual networks implicitly explores the trade-off in allocating parameter resources to width versus depth zagoruyko2016wide, while our focus is on the dynamic allocation of computational resources during inference.
@@ 3.1.2 Modeling Utility and Cost
Computational Cost: Every action taken by an agent incurs a computational cost, $C_{\text{comp}}$. This cost is a function of the resources consumed. For an attention head, the cost could be proportional to the number of tokens it strongly attends to. For an FFN, it could be the number of activated neurons. This aligns with the goals of dynamic computation methods graves2016adaptive, xin2020deebert but provides a more granular, agent-centric view. For a given Transformer layer $l$, we can define its computational cost as a function of its activation patterns. For instance, we can define the cost as the sum of the L1 norm of the attention scores and the FFN activations:
@@ 3.2.1 Implementing Resource Constraints
We implement resource constraints using techniques that induce sparsity in the model’s activations during inference. We focus on constraining the attention mechanism, as it is a primary driver of computational cost and has been a focus of efficiency research kitaev2020reformer, beltagy2020longformer. We employ two main techniques: